# Deep neural networks for choice analysis: Enhancing behavioral regularity with gradient regularization

### v1

@@ Conclusions
DNNs often present behaviorally irregular patterns that greatly limit their practical use and theoretical appeal in travel behavior analysis, especially in applications and forecasting. However, there is no consensus on how to measure or improve the model regularity of DNNs within the field of discrete choice modeling. This paper makes contributions by developing the behavioral regularity metrics and a gradient regularization framework. Specifically, we propose the “law of demand” in economics as a novel measure of DNNs’ behavioral regularity w.r.t. generalized costs. Using a constrained optimization framework, we design six gradient regularizers to enhance the strong and weak behavioral regularities of DNNs. Empirically, these gradient regularizers are applied to two travel survey datasets collected from Chicago and London, through which we examine the trade-off between predictive power and behavioral regularity in the small versus large sample scenarios, and in-domain versus out-of-domain generalizations. Using five evaluation metrics, we demonstrate the effectiveness of our gradient regularizers on both DNNs and TasteNets.
@@
We find that sum-based gradient regularization can significantly improve the behavioral regularity of DNNs without sacrificing their predictive power in all the scenarios. There exists a substitution effect between predictive power and behavioral regularity in the large sample scenario, but a complementary effect in the small sample scenario. This is consistent with general understanding that, within the overparameterized regime, regularization can reduce variance (e.g., Alpaydin, 2014). Utilizing the 10K-Sorted sub-datasets, we further find that gradient regularization is also effective for out-of-domain generalization, which is critical for transferring knowledge across contexts. Our results also demonstrate how and why the gradient regularizers enhance predictive power and behavioral regularity, particularly for the small sample scenario and out-of-domain generalization. Besides the constrained optimization interpretation, the findings could be further understood by an analogy between gradient regularization and informative Bayesian prior. Specifically, gradient regularization can be seen as a deterministic prior imposed on the parameters estimates, that allows incorporating the modeler’s prior belief on the “proper sign” of parameter estimates. Such prior belief can be critical especially when the sample size is limited.
@@
This study pioneers in proposing new behavioral metrics and designing a practical regularization framework for DNNs. To address behavioral irregularity as shown in past studies (Wang et al., 2020a, b; Wong and Farooq, 2021; Xia et al., 2023), we incorporate domain-knowledge into DNNs by regularizing the gradients’ direction, in contrast to the norm-based (magnitude) regularization as in the computer science literature (Drucker and Le Cun, 1991; Jakubovitz and Giryes, 2018; Sokolić et al., 2017). With this research, strong and weak behavioral regularities could be incorporated into future behavioral analysis using deep learning, thus evaluating the consistency of models with theories in behavioral science and microeconomics. Gradient regularization is a general framework, which can impose any prior belief of the input–output relationship to model training. With more behaviorally regular models, future researchers can integrate the computational power of deep learning and the theoretical rigor in behavioral science, thus developing reliable deep learning models for transportation policy analysis. For example, researchers can use our approach to facilitate cross-city transfer of mobility policies when local governments seek to learn certain policies (e.g., congestion charging) from other areas. Our study demonstrates that deep learning with behavioral regularity could generate reliable insights for such policy transfer even when the source and the target cities have different data distributions in sociodemographics and pricing strategies.
@@
Limitations still exist in this study. This work only regularizes direct partial derivatives, but not cross partial derivatives. Regularization on cross partial derivatives, on the other hand, could impose a stronger prior assumption, controlling the substitution and complementary patterns across alternatives. In addition, this research assumes exogenous penalty weight for gradient regularization, future research could combine hyperparameter tuning to automatically learn these penalty weights. Lastly, as in typical deep learning models, parameter identification has been discussed (Hwang and Ding, 1997) but largely remains an open question. In traditional DCMs, parameter identification is the foundation for statistical analysis (McFadden, 1980). However, without further research into DNNs’ parameter identification, it could be challenging to quantify model uncertainty or design statistical tests, thus limiting the practicality of deep learning for discrete choice analysis. Future studies should investigate the necessary and sufficient conditions for model identification and statistical tests, especially in the over-parameterized deep learning models.
Contributions of authors
Shenhao Wang conceived of the presented idea; Shenhao Wang and Rui Yao developed the theory and reviewed previous studies; Shenhao Wang, Rui Yao, and Siqi Feng designed the experiments and discussed the results; Siqi Feng conducted the experiments and drafted the manuscript; Shenhao Wang and Rui Yao revised the manuscript; Stephane Hess, Ricardo Daziano, Timothy Brathwaite, and Joan Walker provided comments, Shenhao Wang supervised this work. All authors contributed to the final manuscript.
Acknowledgement
@@
Stephane Hess acknowledges support from the European Research Council through the advanced grant 101020940-SYNERGY.
@@ Appendix A Summary statistics of datasets
Table A.1: Summary statistics of the full Chicago dataset.
Mean Std. Min. 25% 50% 75% Max.
@@
Gender 11,821 (1: male); 14,278 (0: female)
Bachelor’s degree or above 18,853 (1: yes); 7,246 (0: no)
One-person household 6,697 (1: yes); 19,402 (0: no)
@@
[…]
Table A.2: Summary statistics of the full London dataset.
Mean Std. Min. 25% 50% 75% Max.
@@
Number of transfers in public transit $0.37$ $0.62$ 00 00 00 $1$ $4$
@@
Age (year) 18,917 (1: $<25$); 44,957 (0: 25–55); 17,212 (2: $>55$)
Gender 38,396 (1: male); 42,690 (0: female)
@@
10K-Sorted sub-datasets
Table A.3: Summary statistics of 10K-Sorted (Chicago).
Training set (7K samples)
Mean Std. Min. 25% 50% 75% Max.
@@
Gender 3,104 (1: male); 3,896 (0: female)
Bachelor’s degree or above 5,087 (1: yes); 1,913 (0: no)
One-person household 1,830 (1: yes); 5,170 (0: no)
One-car household 2,678 (1: yes); 4,322 (0: no)
[…]
Test set (2K samples)
@@
Gender 963 (1: male); 1,037 (0: female)
Bachelor’s degree or above 1,439 (1: yes); 561 (0: no)
One-person household 436 (1: yes); 1,564 (0: no)
One-car household 669 (1: yes); 1,331 (0: no)
[…]
Table A.4: Summary statistics of 10K-Sorted (London).
Training set (7K samples)
Mean Std. Min. 25% 50% 75% Max.
@@
Number of transfers in public transit $0.16$ $0.43$ 00 00 00 00 $3$
@@
Age (year) 1,746 (1: $<25$); 3,525 (0: 25–55); 1,729 (2: $>55$)
Gender 3,296 (1: male); 3,704 (0: female)
Driving license 4,140 (1: yes); 2,860 (0: no)
Test set (2K samples)
@@
Number of transfers in public transit $1.20$ $0.57$ 00 $1$ $1$ $1$ $4$
@@
Age (year) 212 (1: $<25$); 1,636 (0: 25–55); 152 (2: $>55$)
Gender 1,028 (1: male); 972 (0: female)
Driving license 1,561 (1: yes); 439 (0: no)
@@ Appendix B Impacts of optimization algorithms
Using the 10K-Random Chicago dataset, we visualize the training and validation losses (Fig. B.1) as well as the corresponding individual demand functions (Fig. B.2) for the Adam, AdamW, and SGD algorithms. Based on default settings of all three algorithms, both Adam and AdamW lead to fast convergence and reasonable demand functions, whereas SGD does not converge within 100 epochs and leads to less reasonable demand functions. It is worth noting that weight decay of AdamW is set as zero to avoid further complexity in regularization, thus identifying the direct impacts of our gradient regularization on behavioral regularity.
@@
Figure B.1: Training loss (blue) and validation loss (orange) per epoch for different algorithms.
@@
Figure B.2: Individual demands as functions of driving cost for different algorithms.
@@ Appendix C Training and validation performance
Table C.1: Model performance in the training sets (10K-Random).
Panel 1: Chicago dataset, sum-XGR
@@
mean (std.) No GR PGR UGR LGR No GR PGR UGR LGR MNL
Log-likelihood $-4272.7$ $\mathbf{-4216.5}$ $-4240.1$ $-4221.9$ $-4677.8$ $-4675.8$ $-4676.0$ $\underline{-4675.7}$ $-4710.2$
(12.97) (11.74) (15.40) (17.58) (19.89) (19.38) (19.57) (19.24) (0)
Accuracy 0.751 0.753 0.752 0.753 0.732 0.732 0.733 0.732 0.731
(0.002) (0.002) (0.002) (0.002) (0.002) (0.002) (0.002) (0.002) (0)
$F_{1}$ score 0.717 0.726 0.721 0.724 0.675 0.675 0.675 0.675 0.681
(0.004) (0.004) (0.003) (0.004) (0.004) (0.004) (0.004) (0.004) (0)
Strong regularity 0.889 0.987 0.980 0.988 0.997 0.998 0.998 0.998 0.998
(0.061) (0.003) (0.011) (0.003) (0.003) (0.002) (0.002) (0.002) (0)
Weak regularity 0.927 0.999 0.997 0.999 0.999 0.999 0.999 0.999 1.000
(0.058) (0.001) (0.004) (0.001) (0.002) (0.001) (0.002) (0.001) (0)
Panel 2: Chicago dataset, norm-XGR
Log-likelihood $\mathbf{-4272.7}$ $-4282.9$ $-4333.2$ $-4303.7$ $\underline{-4677.8}$ $-4679.7$ $-4763.8$ $-4684.8$ $-4710.2$
(13.22) (12.23) (14.33) (11.96) (19.89) (19.88) (17.51) (19.68) (0)
Accuracy 0.751 0.750 0.745 0.748 0.732 0.732 0.729 0.732 0.731
(0.002) (0.002) (0.002) (0.001) (0.002) (0.002) (0.002) (0.002) (0)
$F_{1}$ score 0.717 0.713 0.700 0.706 0.675 0.674 0.665 0.673 0.681
(0.004) (0.004) (0.005) (0.005) (0.004) (0.004) (0.004) (0.004) (0)
Strong regularity 0.889 0.856 0.704 0.813 0.997 0.997 0.922 0.996 0.998
(0.061) (0.063) (0.044) (0.063) (0.003) (0.003) (0.032) (0.005) (0)
Weak regularity 0.927 0.898 0.761 0.855 0.999 0.999 0.935 0.998 1.000
(0.058) (0.062) (0.046) (0.065) (0.002) (0.002) (0.028) (0.003) (0)
Table C.1 (continued).
Panel 3: London dataset, sum-XGR
@@
mean (std.) No GR PGR UGR LGR No GR PGR UGR LGR MNL
Log-likelihood $-4427.9$ $\mathbf{-4416.1}$ $-4434.8$ $-4499.3$ $\underline{-4500.2}$ $-4517.9$ $-4567.1$ $-4518.7$ $-4834.7$
(20.61) (39.09) (24.41) (16.00) (10.52) (12.03) (24.14) (11.93) (0)
Accuracy 0.732 0.729 0.731 0.729 0.736 0.734 0.728 0.734 0.718
(0.003) (0.004) (0.005) (0.004) (0.001) (0.002) (0.004) (0.002) (0)
$F_{1}$ score 0.731 0.728 0.730 0.726 0.732 0.730 0.723 0.730 0.714
(0.002) (0.003) (0.004) (0.004) (0.001) (0.002) (0.004) (0.002) (0)
Strong regularity 0.947 0.993 0.994 0.997 0.936 0.964 0.988 0.975 0.996
(0.032) (0.004) (0.009) (0.003) (0.023) (0.012) (0.009) (0.011) (0)
Weak regularity 0.968 0.999 0.998 1.000 0.960 1.000 1.000 1.000 1.000
(0.025) (0.001) (0.005) (0.000) (0.019) (0.000) (0.000) (0.000) (0)
Table C.2: Model performance in the validation sets (10K-Random).
Panel 1: Chicago dataset, sum-XGR

### v2

@@ Conclusions
DNNs often present behaviorally irregular patterns that greatly limit their practical use and theoretical appeal in travel behavior analysis, especially in applications and forecasting. Nevertheless, there is no consensus on how to measure or improve the model regularity of DNNs within the field of discrete choice modeling. This paper makes contributions by developing the behavioral regularity metrics and a gradient regularization framework. Specifically, we propose the law of demand in economics as a novel measure of DNNs’ behavioral regularity w.r.t. generalized costs. Using a constrained optimization framework, we design six gradient regularizers to enhance the strong and weak behavioral regularities of DNNs. Empirically, these gradient regularizers are applied to two travel survey datasets collected from Chicago and London, through which we examine the trade-off between predictive power and behavioral regularity in the small versus large sample scenarios, and in-domain versus out-of-domain generalizations. Using five evaluation metrics, we demonstrate the effectiveness of our gradient regularizers on both DNNs and TasteNets.
@@
We find that sum-based gradient regularization can significantly improve the behavioral regularity of DNNs without sacrificing their predictive power in all the scenarios. There exists a substitution effect between predictive power and behavioral regularity in the large sample scenario, but a complementary effect in the small sample scenario. This is consistent with general understanding that, within the overparameterized regime, regularization can reduce variance (e.g., Alpaydin, 2014). [^4] To be precise, this coincides with the classical viewpoint of bias–variance trade-off. Recent studies have found evidence that over-parameterization with large data samples could potentially improve model performance (e.g., Belkin et al., 2019). Utilizing the 10K-Sorted datasets, we further find that gradient regularization is also effective for out-of-domain generalization, which is critical for transferring knowledge across contexts. Our results also demonstrate how and why the gradient regularizers enhance predictive power and behavioral regularity, particularly for the small sample scenario and out-of-domain generalization. Comparing the soft and hard constraints, the DNNs with soft constraints can outperform the TasteNets with either soft or hard constraints, implying that TasteNets might be overly restrictive in capturing behavioral mechanisms. Besides the constrained optimization interpretation, the findings could be further understood by an analogy between gradient regularization and informative Bayesian prior. Specifically, gradient regularization can be seen as a deterministic prior imposed on the parameters estimates, that allows incorporating the modeler’s prior belief. Such prior belief can be critical especially when the sample size is limited.
@@
This study pioneers in proposing new behavioral metrics and designing a practical regularization framework for DNNs. To address behavioral irregularity as shown in past studies (Wang et al., 2020a, b; Wong and Farooq, 2021; Xia et al., 2023), we incorporate domain-knowledge into DNNs by regularizing the gradients’ direction, in contrast to the norm-based (magnitude) regularization as in the computer science literature (Drucker and Le Cun, 1991; Jakubovitz and Giryes, 2018; Sokolić et al., 2017). With this research, strong and weak behavioral regularities could be incorporated into future behavioral analysis using deep learning, thus evaluating the consistency of models with theories in behavioral science and microeconomics. Gradient regularization is a practical framework, which can impose any prior belief of the input–output relationship on any NN architecture. The gradient regularization method provides an alternative approach to enhance behavioral regularity, in addition to designing dedicated network architectures. It integrates the computational power of deep learning with common behavioral regularity assumptions through simple gradient regularization. Hence, it provides an opportunity to develop more plausible deep learning models for transportation policy analysis. For example, researchers can use our approach to facilitate cross-city transfer of mobility policies when local governments seek to learn certain policies (e.g., congestion charging) from other areas. Our study demonstrates that deep learning with behavioral regularity could generate reliable insights for such policy transfer even when the source and the target cities have different data distributions in sociodemographics and pricing strategies.
This study mainly adopts the market-level aggregate metric for behavioral regularity and enhances it through soft constraints, despite other possible approaches. Instead of the aggregate regularity, researchers may consider individual behavioral regularity and enhance it through hard constraints (Kim and Bansal, 2024). We focus on the aggregate regularity because the metric is weaker than the individual one, thereby accommodating individual-level irrational behaviors (Bagwell and Riordan, 1991; Stiving, 2000). Empirically, the aggregate regularity and soft constraints jointly lead to higher predictive performance, implying the existence of irregular behaviors for certain individuals or price ranges. In fact, the two issues – individual versus aggregate regularity and soft versus hard constraints – have been debated for a long time in the economics and optimization fields (Becker, 1962), and we avoid taking a strong stance on either side. Nevertheless, the aggregate regularity and soft constraints form the weakest assumption among all possibilities, thus serving as an acceptable common ground for the research community.
@@
Limitations still exist in this study. This work only regularizes direct partial derivatives, but not cross partial derivatives. Regularization on cross partial derivatives, on the other hand, could impose a stronger prior assumption, controlling the substitution and complementary patterns across alternatives. In addition, this research assumes exogenous penalty weight for gradient regularization, while future research could combine hyperparameter tuning to automatically learn these penalty weights. Lastly, as in typical deep learning models, parameter identification has been discussed (Hwang and Ding, 1997) but largely remains an open question. In traditional DCMs, parameter identification is the foundation for statistical analysis (McFadden, 1980). However, without further research into DNNs’ parameter identification, it could be challenging to quantify model uncertainty or design statistical tests, thus limiting the practicality of deep learning for discrete choice analysis. Future studies should investigate the necessary and sufficient conditions for model identification and statistical tests, especially in the over-parameterized deep learning models.
CRediT authorship contribution statement
Siqi Feng: Writing – review & editing, Writing – original draft, Visualization, Validation, Methodology, Investigation, Formal analysis, Data curation. Rui Yao: Writing – review & editing, Writing – original draft, Validation, Methodology, Investigation, Conceptualization. Stephane Hess: Writing – review & editing. Ricardo A. Daziano: Writing – review & editing. Timothy Brathwaite: Writing – review & editing. Joan Walker: Writing – review & editing. Shenhao Wang: Writing – review & editing, Writing – original draft, Validation, Supervision, Project administration, Methodology, Investigation, Formal analysis, Data curation, Conceptualization.
Acknowledgments
@@
We thank the three anonymous reviewers for their constructive criticism and valuable suggestions. We also thank Qingyi Wang, Yunhan Zheng, and Dingyi Zhuang from MIT for their helpful comments. Stephane Hess acknowledges support from the European Research Council through the advanced grant 101020940-SYNERGY.
@@ Appendix A Summary statistics of datasets
See Tables 5 and 6.
Table 5: Summary statistics of the full CMAP dataset.
Variable Mean SD Min. 25% 50% 75% Max.
@@
Male 11,821 (1: yes); 14,278 (0: no)
Bachelor’s or graduate degree 18,853 (1: yes); 7246 (0: no)
One-person household 6697 (1: yes); 19,402 (0: no)
@@
Household income $\geq$ \$75K 15,520 (1: yes); 10,579 (0: no)
Table 6: Summary statistics of the full LTDS dataset.
Variable Mean SD Min. 25% 50% 75% Max.
@@
Number of public transit transfers $0.37$ $0.62$ 00 00 00 $1$ $4$
@@
Youth (age $<$ 25) 18,917 (1: yes); 62,169 (0: no)
Senior (age $>$ 55) 17,212 (1: yes); 63,874 (0: no)
Male 38,396 (1: yes); 42,690 (0: no)
@@
10K-Sorted datasets
See Tables 7, 8, 9 and 10.
Table 7: Summary statistics of the training set of 10K-Sorted CMAP.
Variable Mean SD Min. 25% 50% 75% Max.
@@
Male 3104 (1: yes); 3896 (0: no)
Bachelor’s or graduate degree 5087 (1: yes); 1913 (0: no)
One-person household 1830 (1: yes); 5170 (0: no)
One-car household 2678 (1: yes); 4322 (0: no)
Household income $\geq$ \$75K 4079 (1: yes); 2921 (0: no)
Table 8: Summary statistics of the test set of 10K-Sorted CMAP.
Variable Mean SD Min. 25% 50% 75% Max.
@@
Male 963 (1: yes); 1037 (0: no)
Bachelor’s or graduate degree 1439 (1: yes); 561 (0: no)
One-person household 436 (1: yes); 1564 (0: no)
One-car household 669 (1: yes); 1331 (0: no)
Household income $\geq$ \$75K 1309 (1: yes; 691 (0: no)
Table 9: Summary statistics of the training set of 10K-Sorted LTDS.
Variable Mean SD Min. 25% 50% 75% Max.
@@
Number of public transit transfers $0.16$ $0.43$ 00 00 00 00 $3$
@@
Youth (age $<$ 25) 1746 (1: yes); 5254 (0: no)
Senior (age $>$ 55) 1729 (1: yes); 5271 (0: no)
Male 3296 (1: yes); 3704 (0: no)
Driving license 4140 (1: yes); 2860 (0: no)
Table 10: Summary statistics of the test set of 10K-Sorted LTDS.
Variable Mean SD Min. 25% 50% 75% Max.
@@
Number of public transit transfers $1.20$ $0.57$ 00 $1$ $1$ $1$ $4$
@@
Youth (age $<$ 25) 212 (1: yes); 1788 (0: no)
Senior (age $>$ 55) 152 (1: yes); 1848 (0: no)
Male 1028 (1: yes); 972 (0: no)
Driving license 1561 (1: yes); 439 (0: no)
@@ Appendix B Impacts of optimization algorithms
Using the 10K-Random CMAP dataset, we visualize the training and validation losses (Fig. 8) as well as the corresponding individual demand functions (Fig. 9) for the Adam, AdamW, and SGD algorithms. Based on default settings of all three algorithms, both Adam and AdamW lead to fast convergence and reasonable demand functions, whereas SGD does not converge within 100 epochs and leads to less reasonable demand functions. It is worth noting that weight decay of AdamW is set as zero to avoid further complexity in regularization, thus identifying the direct impacts of our gradient regularization on behavioral regularity.
@@
Figure 8: Training loss (blue) and validation loss (orange) per epoch for different algorithms.
@@
Figure 9: Individual demands as functions of driving costs for different algorithms.
@@ Appendix C Training and validation performance
See Tables 11 and 12.
Table 11: Model performance in the training sets of 10K-Random.
@@
Mean (SD) No GR PGR UGR LGR No GR PGR UGR LGR ReLU Exp MNL
Panel 1: CMAP data, sum-XGR
Log-likelihood $-4272.7$ $\mathbf{-4216.5}$ $-4240.1$ $\underline{-4221.9}$ $-4677.8$ $-4675.8$ $-4676.0$ $-4675.7$ $-4737.2$ $-4675.7$ $-4710.2$
(12.97) (11.74) (15.40) (17.58) (19.89) (19.38) (19.57) (19.24) (51.18) (18.97) (0)
Accuracy 0.751 0.753 0.752 0.753 0.732 0.732 0.733 0.732 0.729 0.733 0.731
(0.002) (0.002) (0.002) (0.002) (0.002) (0.002) (0.002) (0.002) (0.002) (0.002) (0)
$F_{1}$ score 0.717 0.726 0.721 0.724 0.675 0.675 0.675 0.675 0.671 0.681 0.681
(0.004) (0.004) (0.003) (0.004) (0.004) (0.004) (0.004) (0.004) (0.004) (0.003) (0)
Strong regularity 0.889 0.987 0.980 0.988 0.997 0.998 0.998 0.998 0.423 0.999 0.998
(0.061) (0.003) (0.011) (0.003) (0.003) (0.002) (0.002) (0.002) (0.315) (0.000) (0)
Weak regularity 0.927 0.999 0.997 0.999 0.999 0.999 0.999 0.999 1.000 1.000 1.000
(0.058) (0.001) (0.004) (0.001) (0.002) (0.001) (0.002) (0.001) (0.000) (0.000) (0)
Panel 2: CMAP data, norm-XGR
Log-likelihood $\mathbf{-4272.7}$ $\underline{-4282.9}$ $-4333.2$ $-4303.7$ $-4677.8$ $-4679.7$ $-4763.8$ $-4684.8$ $-4737.2$ $-4675.7$ $-4710.2$
(13.22) (12.23) (14.33) (11.96) (19.89) (19.88) (17.51) (19.68) (51.18) (18.97) (0)
Accuracy 0.751 0.750 0.745 0.748 0.732 0.732 0.729 0.732 0.729 0.733 0.731
(0.002) (0.002) (0.002) (0.001) (0.002) (0.002) (0.002) (0.002) (0.002) (0.002) (0)
$F_{1}$ score 0.717 0.713 0.700 0.706 0.675 0.674 0.665 0.673 0.671 0.681 0.681
(0.004) (0.004) (0.005) (0.005) (0.004) (0.004) (0.004) (0.004) (0.004) (0.003) (0)
Strong regularity 0.889 0.856 0.704 0.813 0.997 0.997 0.922 0.996 0.423 0.999 0.998
(0.061) (0.063) (0.044) (0.063) (0.003) (0.003) (0.032) (0.005) (0.315) (0.000) (0)
Weak regularity 0.927 0.898 0.761 0.855 0.999 0.999 0.935 0.998 1.000 1.000 1.000
(0.058) (0.062) (0.046) (0.065) (0.002) (0.002) (0.028) (0.003) 1.000 1.000 (0)
Panel 3: LTDS data, sum-XGR
Log-likelihood $\underline{-4427.9}$ $\mathbf{-4416.1}$ $-4434.8$ $-4499.3$ $-4500.2$ $-4517.9$ $-4567.1$ $-4518.7$ $-4589.4$ $-4498.1$ $-4834.7$
(20.61) (39.09) (24.41) (16.00) (10.52) (12.03) (24.14) (11.93) (68.97) (9.146) (0)
Accuracy 0.732 0.729 0.731 0.729 0.736 0.734 0.728 0.734 0.728 0.738 0.718
(0.003) (0.004) (0.005) (0.004) (0.001) (0.002) (0.004) (0.002) (0.005) (0.001) (0)
$F_{1}$ score 0.731 0.728 0.730 0.726 0.732 0.730 0.723 0.730 0.724 0.734 0.714
(0.002) (0.003) (0.004) (0.004) (0.001) (0.002) (0.004) (0.002) (0.005) (0.001) (0)
Strong regularity 0.947 0.993 0.994 0.997 0.936 0.964 0.988 0.975 0.927 0.990 0.996
(0.032) (0.004) (0.009) (0.003) (0.023) (0.012) (0.009) (0.011) (0.046) (0.002) (0)
Weak regularity 0.968 0.999 0.998 1.000 0.960 1.000 1.000 1.000 1.000 1.000 1.000
(0.025) (0.001) (0.005) (0.000) (0.019) (0.000) (0.000) (0.000) (0.000) (0.000) (0)
Table 12: Model performance in the validation sets of 10K-Random.
@@
Mean (SD) No GR PGR UGR LGR No GR PGR UGR LGR ReLU Exp MNL
Panel 1: CMAP data, sum-XGR
Log-likelihood $\underline{-633.6}$ $\mathbf{-633.0}$ $-634.1$ $-634.2$ $-683.0$ $-683.1$ $-683.1$ $-683.1$ $-689.2$ $-687.1$ $-693.8$
(2.507) (2.474) (3.594) (2.794) (2.442) (2.470) (2.482) (2.467) (6.598) (2.722) (0)
Accuracy 0.739 0.739 0.739 0.739 0.726 0.727 0.727 0.727 0.724 0.728 0.737
(0.004) (0.004) (0.003) (0.003) (0.004) (0.003) (0.003) (0.003) (0.003) (0.004) (0)
$F_{1}$ score 0.702 0.710 0.705 0.708 0.668 0.669 0.670 0.669 0.664 0.679 0.690
(0.005) (0.004) (0.005) (0.003) (0.007) (0.006) (0.006) (0.006) (0.003) (0.004) (0)
Strong regularity 0.879 0.981 0.973 0.982 0.997 0.999 0.998 0.998 0.427 1.000 1.000
(0.058) (0.005) (0.014) (0.005) (0.005) (0.002) (0.003) (0.002) (0.319) (0.001) (0)
Weak regularity 0.928 0.999 0.997 0.999 0.998 0.999 0.999 0.999 1.000 1.000 1.000
(0.057) (0.001) (0.004) (0.001) (0.003) (0.002) (0.002) (0.002) (0.000) (0.000) (0)
Panel 2: CMAP data, norm-XGR
Log-likelihood $\underline{-633.6}$ $\mathbf{-633.3}$ $-635.9$ $\mathbf{-633.3}$ $-683.0$ $-683.0$ $-687.4$ $-682.9$ $-689.2$ $-687.1$ $-693.8$
(2.507) (2.414) (2.676) (2.989) (2.442) (2.421) (2.236) (2.366) (6.598) (2.722) (0)
Accuracy 0.739 0.741 0.741 0.741 0.726 0.725 0.720 0.725 0.724 0.728 0.737
(0.004) (0.004) (0.003) (0.003) (0.004) (0.004) (0.001) (0.004) (0.003) (0.004) (0)
$F_{1}$ score 0.702 0.702 0.696 0.699 0.668 0.666 0.655 0.665 0.664 0.679 0.690
(0.005) (0.005) (0.006) (0.005) (0.007) (0.007) (0.003) (0.006) (0.003) (0.004) (0)
Strong regularity 0.879 0.846 0.701 0.803 0.997 0.996 0.924 0.995 0.427 1.000 1.000
(0.058) (0.058) (0.044) (0.061) (0.005) (0.005) (0.03) (0.007) (0.319) (0.001) (0)
Weak regularity 0.928 0.900 0.766 0.858 0.998 0.998 0.935 0.997 1.000 1.000 1.000
(0.057) (0.057) (0.046) (0.062) (0.003) (0.003) (0.026) (0.006) (0.000) (0.000) (0)
Panel 3: LTDS data, sum-XGR
Log-likelihood $-645.5$ $-642.8$ $-644.2$ $-647.8$ $\mathbf{-638.8}$ $\underline{-639.9}$ $-645.3$ $\underline{-639.9}$ $-654.1$ $-640.6$ $-670.0$
(4.399) (5.709) (4.345) (3.915) (1.949) (2.182) (3.254) (1.985) (11.75) (2.157) (0)
Accuracy 0.721 0.722 0.725 0.723 0.729 0.726 0.726 0.726 0.720 0.726 0.735
(0.002) (0.003) (0.002) (0.005) (0.003) (0.002) (0.002) (0.002) (0.006) (0.003) (0)
$F_{1}$ score 0.719 0.721 0.722 0.718 0.723 0.720 0.719 0.720 0.714 0.721 0.730
(0.003) (0.003) (0.002) (0.004) (0.003) (0.002) (0.003) (0.002) (0.007) (0.003) (0)
Strong regularity 0.947 0.992 0.991 0.996 0.924 0.953 0.982 0.964 0.916 0.982 0.991
(0.032) (0.005) (0.012) (0.004) (0.026) (0.014) (0.011) (0.014) (0.050) (0.003) (0)
Weak regularity 0.971 0.999 0.997 1.000 0.957 1.000 1.000 1.000 1.000 1.000 1.000
(0.025) (0.001) (0.007) (0.001) (0.023) (0.001) (0.000) (0.001) (0.000) (0.000) (0)
Small sample scenario
See Tables 13 and 14.
Table 13: Model performance in the training sets of 1K-Random.